If jenna cleans her house in 4 hours and jodi,her daughter,cleans the house in 6 hours, how long will it talke for them to clean the house together?
Answer by Cromlix(4381)

Hi there
Jenna can clean her house in 4hours
Jodi can clean her house in 6hours
1/4 + 1/6 = 1/x
Common denominator = 24x
6x + 4x = 24
10x = 24
x = 2.4 hours
x = 2 hours 24 mins
This is how long they take together to clean
the house.
Hope this helps:-)